OpenStudy (anonymous):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

Directrix (directrix):

Closed over multiplication in this problem would mean that if you computer 0*0, 0*1, or 1*1, then you would always get back as the answer one of the two elements in the set, that is you'd always get 0 or1. So, if you do all those multiplications, do you always get 0 or 1 back? Then, you have to test division next.

OpenStudy (anonymous):

so a and d

OpenStudy (michele_laino):

@ineedurhelpnow under what topology?

Directrix (directrix):

On subtraction, then what would happen if you did the calculation 0 - 1? Would you get 0 or 1 as the answer?

Directrix (directrix):

@ineedurhelpnow You need to reconsider you answer of d.

OpenStudy (anonymous):

ok then what do you think it would be

Directrix (directrix):

First, what is the number zero with one subtracted from it equal to? Is it 0 or 1 ?

OpenStudy (anonymous):

-1

Directrix (directrix):

Yes. Is negative one a member of the original set?

Directrix (directrix):

So is the set closed under subtraction?

OpenStudy (anonymous):

uhhhhhh i think

Directrix (directrix):

It is* not closed* under subtraction because we took two elements of the set, subtracted one from the other, and got an answer that was not in the set.

Directrix (directrix):

That leaves just A as the answer.

Directrix (directrix):

Question?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

First, close this thread and then post the question in a new thread. Once you do that, I'm going to ask if you know how the set of whole numbers looks.

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

a. 0 + (-1)=-1 d.2+(-1)=1

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

ok can u just give me the answer this test is timed

Directrix (directrix):

WHOLE NUMBER Definition of Whole Numbers The numbers in the set {0, 1, 2, 3, 4, 5, 6, 7, . . . . } are called whole numbers. In other words, whole numbers is the set of all counting numbers plus zero. http://www.icoachmath.com/math_dictionary/whole_numbers.html

OpenStudy (anonymous):

please just tell me i know what a whole number is

Directrix (directrix):

Whole numbers are not negative. So, both of these are WRONG.

OpenStudy (anonymous):

so its b,c,d

OpenStudy (anonymous):

ok i got 45 more sec on this question

OpenStudy (anonymous):

tellllllllllllllll meeeeeeeeeeeeeeeeee

Directrix (directrix):

A is wrong because -1 is not a whole number.

OpenStudy (anonymous):

ok i have 5 min on this question Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2} can u just give me the answer that would help

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

HELLLLLLLLLLLLLLLLLLLLLLLLLLLLLP

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

I cannot feed you answers on this site because it is against the Code of Conduct.

OpenStudy (anonymous):

well HELPPPPPPP and ppl do it all the time

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok

Directrix (directrix):

The rational numbers are closed under subtraction. There is no easy way to show that. B is correct.

Directrix (directrix):

What about this: C. {0, 1} What is the answer to 0 - 1 ? Is that answer a member of the given set?

OpenStudy (anonymous):

-1

Directrix (directrix):

So, no. C is not the correct answer.

OpenStudy (anonymous):

so it is d and b

Directrix (directrix):

So, far. we have that B is the only correct answer. But, we are not finished.

OpenStudy (anonymous):

Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct. A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. is a positive integer

Directrix (directrix):

{0, 1, 2} What is the answer to 1-2 ? Is that a member of the given set?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yasssssss it has to be more than one hun

OpenStudy (anonymous):

yeah could you help with Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct. A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. is a positive intege

OpenStudy (anonymous):

@Directrix